# Semi-SD: Semi-Supervised Metric Depth Estimation via Surrounding Cameras for Autonomous Driving

### v1

@@ IV-B Qualitative and quantitative analysis
In Sec. IV-C, we evaluate the model’s performance across varying levels of complexity. To ensure a fair comparison, we used our medium model (Ours-Medium), which employs the same backbone as the baseline, for benchmarking.
@@
Per-Camera Evaluation. We also do per-camera evaluation in Table II. M2-Depth [15] estimate the pose by the front camera, so the metric  Abs.Sql on front camera is the best. But our framework can achieve a error banlance across surrounding cameras by global feature fusion.
@@ IV-C Ablation study
TABLE III: Qualitative comparison results on DDAD [1] (evaluate depth: 80 m) of the three modules: spatial-temporal attention (ST), SAM feature (SAM), depth-enhanced motion estimation (DM).
@@
TABLE IV: Qualitative comparison results on DDAD [1] (evaluate depth: 80 m) of the two losses: depth loss $\mathcal{L}_{d}$ and curvature loss $\mathcal{L}_{\nabla}$.
@@
Figure 8: Visualization of training iterations on DDAD [1] dataset. It can be observed that as we progressively incorporate the proposed modules, the overall model’s metric accuracy and convergence speed improve steadily. $val/abs\_rel$ and $val/a1$ represent $Abs.Rel$ and $\delta<1.25$, respectively.
@@
Ablation Study on Model Complexity. We conducted experiments using ResNet [18] architectures with varying numbers of layers and transformer structures with different blocks and embedding feature dimensions. Three versions of the implementation are provided, with detailed configurations and quantitative performance metrics on nuScenes dataset [34] presented in Table V. For the larger nuScenes dataset [34], larger models exhibit superior performance, highlighting the importance of dataset scale in model generalization.
TABLE V: The metric performance of different parameter settings on the nuScenes [34] dataset. $n_{l}$ refers to the layer of the main feature extraction ResNet network. $n_{b}$ and $dim_{e}$ represent the number of blocks and the feature dimensions, respectively.
Model $n_{l}$ $n_{b}$ $dim_{e}$ Abs.Rel. Sq.Rel. RMSE RMSE log $\delta<1.25^{3}$
Ours-Small 18 2 16 0.219 2.823 6.238 0.314 0.872
Ours-Medium 34 4 32 0.197 2.624 6.094 0.297 0.903
Ours-Large 50 8 64 0.190 2.523 6.137 0.283 0.912
@@ IV-D GPU Usage and Time Consumption
We compare the GPU usage and time consumption of our models with other baselines, as shown in Table VI. Note that we use the semantic world model [32] to obtain online semantic features, which we have taken into consideration. It is evident that our model demonstrates improvements in time efficiency. Although M2-Depth [15] claims to also utilize semantic features, we cannot verify whether they use the offline semantic world model to reduce the GPU usage due to the lack of source code.
TABLE VI: Comparison of GPU usage and inference time of one frame.
@@
Ours-Small 8902 215
Ours-Medium 9137 232
Ours-Large 9540 239
We also validate the computation time consumed by different modules, as shown in Table VII. The modules with different parameter settings have been underlined, and significant differences on time consumption can be observed.
TABLE VII: The computation time consumed by the different modules, including modules Feature Network (FN), Online Semantic Feature (OSF), Semantic Feature Adoptor (SFA), Spatial-temporal-semantic Transformer (STST), Depth Decoder (DD) and Pose Decoder (PD).
Methods Modules (milliseconds)
FN OSF SFA STST DD PD
Ours-Small 22.971 110.108 18.632 19.319 44.625 1.523
Ours-Medium 25.781 109.283 18.234 33.183 44.309 1.524
Ours-Large 25.772 110.234 18.830 42.545 44.220 1.523
@@ V Conclusion
In this paper, we propose Semi-SD, a unified approach for metric depth prediction and pose estimation through spatial-temporal-semantic information fusion for surrounding cameras. By designing a unified transformer architecture, we effectively integrate these features, improving computational efficiency and reducing boundary ambiguity. We also introduce a joint pose estimation network for surrounding cameras that combines depth predictions with camera extrinsic parameters to achieve accurate scale estimation. Additionally, we propose a curvature loss function, where guidance from the depth estimation world model significantly accelerates convergence and improves depth prediction accuracy. Experimental results on two widely used datasets demonstrate the superiority of our method, showcasing its broad applicability in autonomous driving systems with surrounding cameras.

### v3

@@ IV-B Qualitative and quantitative analysis
Per-Camera Evaluation. We also do per-camera evaluation in Table II. M2-Depth [14] estimate the pose by the front camera, so the metric  Abs.Sql on front camera is the best. But our framework can achieve a error balance across surrounding cameras by global feature fusion.
@@ IV-C Ablation study
TABLE III: Qualitative comparison results on DDAD (evaluate depth: 80 m) of the three modules: spatial-temporal attention (ST), SAM feature (SAM), depth-enhanced motion estimation (DM).
@@
TABLE IV: Qualitative comparison results on DDAD (evaluate depth: 80 m) of the two losses: depth loss $\mathcal{L}_{d}$ and curvature loss $\mathcal{L}_{\nabla}$.
@@
Figure 8: Visualization of training iterations on DDAD [16] dataset. It can be observed that as we progressively incorporate the proposed modules, the overall model’s metric accuracy and convergence speed improve steadily. val/abs_relval/abs\_rel and $val/a1$ represent $Abs.Rel$ and $\delta<1.25$, respectively.
@@ IV-D GPU Usage and Time Consumption
We compare the GPU usage and time consumption of our models with other baselines, as shown in Table V. Note that we use the semantic world model [31] to obtain online semantic features, which we have taken into consideration. It is evident that our model demonstrates improvements in time efficiency. Although M2-Depth [14] claims to also utilize semantic features, we cannot verify whether they use the offline semantic world model to reduce the GPU usage due to the lack of source code.
TABLE V: Comparison of GPU usage and inference time of one frame. Best results are underlined.
@@
Ours 9137 232
@@ V Conclusion
In this paper, we propose Semi-SMD, a unified approach for metric depth prediction and pose estimation through spatial-temporal-semantic information fusion for surrounding cameras. By designing a unified transformer architecture, we effectively integrate these features, improving computational efficiency and reducing boundary ambiguity. We also introduce a joint pose estimation network for surrounding cameras that combines depth predictions with camera extrinsic parameters to achieve accurate scale estimation. Additionally, we propose a curvature loss function, where guidance from the depth estimation world model significantly accelerates convergence and improves depth prediction accuracy. Experimental results on two widely used datasets demonstrate the superiority of our method, showcasing its broad applicability in autonomous driving systems with surrounding cameras.